OpenStudy (hba):

In the xy-plane,the graph of the function h is a line.If h(-1)=4 and h(5)=1,what is the value of h(0) ?

OpenStudy (anonymous):

I think it will b helpful if you consider h as f(x) that is H(-1)=4 means... when value of x is -1 ; y value (f(x)= is 4... and soon.. construct the the line equation using two point form.. then find value for h(0)

OpenStudy (hba):

(-1,4) and (5,1)

OpenStudy (anonymous):

We can find 2 points from h(-1)=4 and h(5)=1

OpenStudy (anonymous):

yes.. you got it

OpenStudy (anonymous):

now construct line equation

OpenStudy (hba):

What next ?

OpenStudy (anonymous):

Use the points to find the slope

OpenStudy (hba):

y-y1=m(x-x1) ??

OpenStudy (anonymous):

yes

OpenStudy (hba):

m=y2-y1/x2-x1

OpenStudy (anonymous):

yes

OpenStudy (hba):

m=-1/2

OpenStudy (hba):

y+1=-1/2(x-4)

OpenStudy (anonymous):

??? what is x1 and y1

OpenStudy (hba):

(x1,y1)=(-1,4)

OpenStudy (anonymous):

y+1=-1/2(x-4) is it y-y1=m(x-x1) ??? can you please check...

OpenStudy (anonymous):

your slope is correct...but the point you took is wrong

OpenStudy (hba):

No it is correct :)

OpenStudy (hba):

Do i have to take the other two points ?

OpenStudy (anonymous):

is it...I have doubt.. y1 is 4 and x1 is-1... but in the abopve equation you took y1 as -1 and x1 as 4 amI right?

OpenStudy (hba):

My bad :(

OpenStudy (anonymous):

:) correct it

OpenStudy (hba):

y-4=-1/2(x+1)

OpenStudy (anonymous):

yes.. now find h(0)

OpenStudy (hba):

Ok so x=0

OpenStudy (anonymous):

yes

OpenStudy (hba):

7/2=y < So 3.5 right ?

OpenStudy (anonymous):

yes.. good.. you got it

OpenStudy (hba):

Thanks alot @Rosh007

OpenStudy (anonymous):

welcome hba